OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Define your variable and write an equation that models the cost of each bracelet.

OpenStudy (anonymous):

I think each bracelet cost 6

OpenStudy (anonymous):

is that right

OpenStudy (texaschic101):

8b = 54 - 6 ---- b stands for cost of each bracelet 8b = 48 b = 48/8 b = 6 yep...bracelets are $6 a piece